IF THERE ARE THREE POSSIBLE VALUES FOR SPIN QUANTUM NUMBER, THEN POSSIBLE NUMBER OF ELEMENTS IN THE 5TH PERIOD OF PERIODIC TABLE WOULD BE-
ANS 27
HOW??
PLEASE REPLY FAST

Grade:12th pass

2 Answers

Ankit Sharma
8 years ago
just some enquiry about the question 
 
is the atomic number is 27 or anything else
n if atomic number is  27 i.e cobalt then it would be in 4th period not 5th  isnt it ??
and any other correction in question u can tell ??
Satvir kumar
6 years ago
I in V Orbit there are 5 best holiday and V P orbital is Orbit or 1 in number and 5 power 3 in number and D orbital 5 in number then total 27 elements are there5s=1*34d=5*35p=3*3So totalno of element is 27Hope this will be true answer.....